OpenStudy (anonymous):

Solve for x: 2/3 (x − 2) = 4x.

OpenStudy (anonymous):

2/3 (x − 2) = 4x We dont like fractions that much, so you can simplify. How?

OpenStudy (anonymous):

how to get rid of 2/3? That is division. So you need the opposite to get rid of.

OpenStudy (anonymous):

multiply by 3 both sides 2/3 (x − 2) *3= 4x*3 2(x-2)=12x

OpenStudy (anonymous):

again you can simplify

OpenStudy (anonymous):

2 is a factor of the left hand side and RHS

OpenStudy (anonymous):

So you divide by 2 2(x-2)=12x (x-2)=6x

OpenStudy (anonymous):

the bracket is not needed anymore x-2=6x Take the same parts to the same side of he equation

OpenStudy (anonymous):

-x both sides -2=5x

OpenStudy (anonymous):

divide by 5 to get x

OpenStudy (anonymous):

hope it is clear

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

cool